# Unsupervised Model Learning for Quantum Impurity Systems

### v1

@@ III Single and double extended effective impurity models
where we have assumed particle hole symmetry and $|\phi,\sigma\rangle\equiv|\phi\rangle_{0}\otimes|\sigma\rangle_{d}$ with $\phi=\{0,\uparrow,\downarrow,\uparrow\downarrow\}$ for the local bath orbital and $\sigma=\{\operatorname{\Uparrow},\operatorname{\Downarrow}\}$ for the impurity spin. The corresponding projectors, that are measured in the bare AIM must include an additional multiplier $\hat{\Omega}_{\rm ad}=\hat{n}^{d}_{\uparrow}+\hat{n}^{d}_{\downarrow}-2\hat{n}^{d}_{\uparrow}\hat{n}^{d}_{\downarrow}$ to project out impurity states with 0 and 2 electrons $\hat{P}_{QS}\to\hat{\Omega}_{\rm ad}\cdot\hat{P}_{QS}\cdot\hat{\Omega}_{\rm ad}$. Thus, the representation of $\hat{P}_{QS}$ in second quantized operators is different for bare and effective model. The effective Hamiltonian $\hat{H}^{1n-\rm mol}_{\text{eff}}=\hat{\tilde{H}}_{\text{bath}}+\theta_{0,1}\hat{P}_{0,1}+\theta_{0,0}\hat{P}_{0,0}+\theta_{\pm 1,1/2}\hat{P}_{\pm 1,1/2}$ reads in second quantized notation
$\hat{P}_{0,1}$ $=\frac{3}{4}(\hat{n}^{0}_{\uparrow}+\hat{n}^{0}_{\downarrow}-2\hat{n}^{0}_{\uparrow}\hat{n}^{1}_{\downarrow})+\hat{\mathbf{S}}_{d}\cdot\hat{\mathbf{S}}_{0}$ (32)
$P_{0,0}$ $=\frac{1}{4}(\hat{n}^{1}_{\uparrow}+\hat{n}^{1}_{\downarrow}-2\hat{n}^{1}_{\uparrow}\hat{n}^{1}_{\downarrow})-\mathbf{S}_{1}\cdot\mathbf{S}_{2}$ (33)
$\hat{P}_{\pm 1,1/2}$ $=\hat{\mathbb{I}}_{\rm eff}-\hat{n}^{0}_{\uparrow}-\hat{n}^{0}_{\downarrow}+2\hat{n}^{0}_{\uparrow}\hat{n}^{0}_{\downarrow}\;,$ (34)
@@
where $\hat{n}_{0}=\sum_{\sigma}\hat{c}_{0\sigma}^{\dagger}\hat{c}_{0\sigma}$. The projectors Eq. 32 combined form the most general SU(2) and particle-hole symmetric Hamiltonian on $\mathcal{H}^{\rm 1e-imp}_{\rm eff}$. Recombining the projectors yields a different parametrization with the same expressibility
@@
$\hat{H}^{JU}_{\text{eff}}=\hat{\tilde{H}}_{\text{bath}}+J\hat{\mathbf{S}}_{d}\cdot\hat{\mathbf{S}}_{0}+\tfrac{1}{2}U_{0}(\hat{n}_{0}-\hat{\mathbb{I}}_{\rm eff})^{2}+L\hat{\mathbb{I}}_{\rm eff}\;.$ (35)
This parametrization Eq. 35 is still meaningful in bare and effective model, because of the simplicity of the AIM. However, note that Eq. 35 includes the identity operator. As discussed before the identity is an unmatchable observable between bare and effective model and since it does not have any impact on the physics we set $L=0$ and find $J=\theta_{0,1},~{}U_{0}=2\theta_{\pm 1,1/2}-\tfrac{3}{2}\theta_{0,1}$. Later in the paper we compare in detail the performance of Eq. 35 and Eq. 32. The parametrization Eq. 35 shows that the $\hat{H}^{1n-\rm mol}_{\text{eff}}$ is equivalent to the LM FP Hamiltonian including the leading irrelevant perturbation, as shown in the RG-Analysis by Krishnamurty et. al. [50]. This exemplifies how the extended impurity approach allows to include irrelevant terms in the Hamiltonian.
We can use the same procedure as for the 1e-imp to derive the 2e-imp minimally constrained Hamiltonian. Constructing the Q and S basis of $\mathcal{H}^{\rm 2e-imp}_{\rm eff}$ one finds a total of eight $Q$ and $S$ combinations, but a total of 15 $s_{z}$-multiplets. Since multiplets are arbitrary they can not be compared between bare and effective model. Thus, we need additional operators that commute with the representations of Q and S such that a common eigenbasis can be formed and then allows us to identify every multiplet. Choosing specific operators creates combinations of multiplets that are not all equally suited to express the physics of the problem, we want all terms in the Hamiltonian to commute there are no terms in the Hamiltonian between QN subspaces. Hence one has to use physical intuition to make a good choice for the additional operators, however if we sacrifice the convexity in parts we can introduce terms that hybridize QN subspaces and thus again yield the most general SU(2) symmetric Hamiltonian. First we consider no such terms though, thus we introduce the hopping operator between the first and second Wilson chain sites
$\hat{T}=\sum_{\sigma}(\hat{c}^{\dagger}_{\sigma 0}\hat{c}_{\sigma 1}+\hat{c}^{\dagger}_{\sigma 1}\hat{c}_{\sigma 0})\;.$ (36)
The $\hat{T}$ operator has the advantage that it is supported on the bath Fock-space and is thus comparable between effective and bare Fock space as it does not depend on the impurity Fock-space. Then there is a clear physical motivation why one would want to include this term, that is because it is also present in the bare Hamiltonian. Forming the common eigenspaces of S, Q and now also $\hat{T}$ we find that only one multiplet remains degenerate
|$\ket{0,\frac{1}{2},0,\frac{1}{2};0}$ $=\sqrt{\frac{2}{3}}\ket{\operatorname{\uparrow},\operatorname{\uparrow},\operatorname{\Downarrow}}-\sqrt{\frac{1}{6}}(\ket{\operatorname{\downarrow},\operatorname{\uparrow},\operatorname{\Uparrow}}+\ket{\operatorname{\uparrow},\operatorname{\downarrow},\operatorname{\Uparrow}})$
|$\ket{0,\frac{1}{2},0,-\frac{1}{2};0}$ $=-\sqrt{\frac{2}{3}}\ket{\operatorname{\downarrow},\operatorname{\downarrow},\operatorname{\Uparrow}}+\sqrt{\frac{1}{6}}(\ket{\operatorname{\downarrow},\operatorname{\uparrow},\operatorname{\Downarrow}}+\ket{\operatorname{\uparrow},\operatorname{\downarrow},\operatorname{\Downarrow}})$
|$\ket{0,\frac{1}{2},0,\frac{1}{2};1}$ $=\frac{1}{\sqrt{2}}(\ket{\operatorname{\uparrow\downarrow},0,\operatorname{\Uparrow}}-\ket{0,\operatorname{\uparrow\downarrow},\operatorname{\Uparrow}})$
|$\ket{0,\frac{1}{2},0,-\frac{1}{2};1}$ $=\frac{1}{\sqrt{2}}(\ket{\operatorname{\uparrow\downarrow},0,\operatorname{\Downarrow}}-\ket{0,\operatorname{\uparrow\downarrow},\operatorname{\Downarrow}})\;,$ (37)
where we have adopted the notation |$\ket{Q,S,T,s_{z};m}$ with $T$ being the hopping eigenvalue. This last ambiguity can be resolved by introducing on more operator
$\hat{W}=\hat{n}^{0}_{\operatorname{\uparrow}}\hat{n}^{0}_{\operatorname{\downarrow}}+\hat{n}^{1}_{\operatorname{\uparrow}}\hat{n}^{1}_{\operatorname{\downarrow}}\;,$ (38)
which we only use in the $S=1/2$ and $Q=0$ subspace. Since, the hopping eigenvalues in this subspace are $T\in\{-2,0,2\}$ we do not introduce an additional quantum number for $\hat{W}$, but combine the $T$ and $W$ quantum numbers to $T\in\{-2,0,1,2\}$ such that |$\ket{0,\frac{1}{2},0;0}\rightarrow\ket{0,\frac{1}{2},0}$ and |$\ket{0,\frac{1}{2},0;1}\rightarrow\ket{0,\frac{1}{2},1}$. All states and quantum numbers for this particular basis are given in Appendix B. With the basis being now complete and unique we have again a projector basis $\hat{P}_{QST}$ that allows us to build the effective Hamiltonian $\hat{H}^{\rm 2e-imp}_{\text{eff}}=\hat{\tilde{H}}_{\text{bath}}+\sum_{\{QST\}:\rm ad}\theta_{QST}\hat{P}_{QST}$ while yielding a convex KLD.
In building the 2e-imp basis we have prioritised the convexity over generality of the Hamiltonian, we can also prioritise generality over convexity. The generality is limited because $Q,S$-multiplets are not connected by any term in the Hamiltonian despite Q and S symmetry in principle allowing such transitions. To get the most general Q- and S-symmetric Hamiltonian we can include the terms
$\hat{P}_{QS;t,t^{\prime}}=|Q,S,T=t\rangle\langle Q,S,T=t^{\prime}|+h.c.\;,$ (39)
that connect states with the same $Q,S$ but different $T$ and $W$ eigenvalues. In the next section we show numerical results for the 1e-imp and 2e-imp, where we also compare the performance of the convexity prioritised and generality prioritised 2e-imp Hamiltonians.
IV Quantifying the learning progress
In the previous section we have introduced the KLD loss function Eq. 8 that quantifies the distinguishability between bare and effective model. In general ML applications, the value of the loss function is a useful metric to judge the training progress and performance of the model. In the case of the KLD for impurity models it is impossible to compute the actual value and thus, the training progress cannot be known. However, to optimize the KLD it is not necessary to evaluate it: as we know its gradient Eq. 9 we can perform a gradient descent (GD) optimization. Since the KLD can be shown to be convex for the minimally constrained models, the gradient vanishes only at saddle-points or global minima of the KLD. Thus, the gradient is itself a good replacement for the loss function to determine whether the training of a model has converged. Due to the special form of the gradient an optimum can only be reached when expectation values associated to the effective Hamiltonian interactions coincide in the bare and effective system

### v2

@@ 0 $1$ $0,~{}\pm 1$
where $|\phi,\sigma\rangle\equiv|\phi\rangle_{0}\otimes|\sigma\rangle_{d}$ with $\phi=\{0,\uparrow,\downarrow,\uparrow\downarrow\}$ for the local bath orbital and $\sigma=\{\operatorname{\Uparrow},\operatorname{\Downarrow}\}$ for the impurity spin.
The identity operator can also be resolved in this basis,
$\hat{\mathbbm{1}}_{\rm eff}=\sum_{Q,S}\hat{P}_{QS}\;,$ (44)
such that $\langle\hat{\mathbbm{1}}_{\rm eff}\rangle_{\rm eff}=1$. We can therefore use the identity operator to eliminate one of the projectors in Eq. 42. As an example we eliminate the $\hat{P}_{0,0}$ projector and write,
$\tilde{H}^{\rm imp}_{\rm eff}(\boldsymbol{\theta})=\theta_{\pm 1,1/2}(\hat{P}_{-1,1/2}+\hat{P}_{+1,1/2})+\theta_{0,1}\hat{P}_{0,1}+\theta_{\mathbbm{1}}\hat{\mathbbm{1}}_{\rm eff}\;,$ (45)
where $\theta_{\mathbbm{1}}=\theta_{0,0}$ and the (as yet still unknown) parameters $\theta_{\pm 1,1/2}$ and $\theta_{0,1}$ have been suitably rescaled. This is useful because the identity operator strictly does not flow under RG. Thus, the term $\theta_{\mathbbm{1}}\hat{\mathbbm{1}}_{\rm eff}$ cannot change the RG flow of the effective model or affect thermal expectation values of the projectors $\langle\hat{P}_{QS}\rangle_{\rm eff}$ evaluated in the effective model. Further, setting $\theta_{\mathbbm{1}}=0$ corresponds to a uniform shift of all energy levels and does not affect the dynamics. For the symmetric AIM, the ground state lives in the $Q=0$, $S=0$ sector, so this choice is equivalent to measuring energies relative to the ground state energy. Importantly, the UML optimization, which is based on matching thermal expectation values, is sensitive to this elimination of a projector, as the unassuming operator $\hat{\mathbbm{1}}_{\rm eff}$ becomes non-trivial in the bare model ${\hat{\Omega}^{\dagger}_{\rm ad}\cdot\hat{\mathbbm{1}}_{\rm eff}\cdot\hat{\Omega}_{\rm ad}}\neq\hat{\mathbbm{1}}_{\rm bare}$. The intricacies that arise from this sensitivity are discussed in the detail in Sec. III.2.
In any case we are left with an MC effective model featuring just two tunable parameters. The physical interpretation of Eq. 45 is manifested by expressing the projectors in second quantized form,
$\hat{P}_{0,1}$ $=\frac{3}{4}(\hat{n}_{0\uparrow}+\hat{n}_{0\downarrow}-2\hat{n}_{0\uparrow}\hat{n}_{0\downarrow})+\hat{\mathbf{S}}_{d}\cdot\hat{\mathbf{S}}_{0}$
$\hat{P}_{0,0}$ $=\frac{1}{4}(\hat{n}_{0\uparrow}+\hat{n}_{0\downarrow}-2\hat{n}_{0\uparrow}\hat{n}_{0\downarrow})-\hat{\mathbf{S}}_{d}\cdot\hat{\mathbf{S}}_{0}$ (46)
$\hat{P}_{\pm 1,1/2}$ $=\hat{\mathbbm{1}}_{\rm eff}-\hat{n}_{0\uparrow}-\hat{n}_{0\downarrow}+2\hat{n}_{0\uparrow}\hat{n}_{0\downarrow}\;,$
@@
where $\hat{n}_{0\sigma}=\hat{c}_{0\sigma}^{\dagger}\hat{c}_{0\sigma}^{\phantom{\dagger}}$. Recombining these projectors yields a different parameterization of Eq. 45 (but with the same expressibility). We write it as,
@@
$\tilde{H}^{\rm imp}_{\rm eff}(\boldsymbol{\theta})\equiv\hat{H}^{JU_{0}}_{\text{eff}}=J\hat{\mathbf{S}}_{d}\cdot\hat{\mathbf{S}}_{0}+\tfrac{1}{2}U_{0}(\hat{n}_{0}-\hat{\mathbbm{1}}_{\rm eff})^{2}+L\hat{\mathbbm{1}}_{\rm eff}\;,$ (47)
where $\hat{n}_{0}=\sum_{\sigma}\hat{n}_{0\sigma}$ and $J=\theta_{0,1}-\theta_{0,0}$, $U_{0}=2\theta_{\pm 1,1/2}-\tfrac{3}{2}\theta_{0,1}-\tfrac{1}{2}\theta_{0,0}$, $L=\tfrac{3}{4}\theta_{0,1}+\tfrac{1}{4}\theta_{0,0}+\theta_{\mathbbm{1}}$. Again, the term involving $\hat{\mathbbm{1}}_{\rm eff}$ can be neglected and we set $L=0$. The effective model written in this way is equivalent to a Kondo model with an additional electron-electron interaction term on the local bath orbital.
Eq. 45, and equivalently Eq. 47, are the most general forms of the effective model that live in the extended Fock space ~$\tilde{\mathcal{H}}^{\rm imp}_{\rm eff}$, consistent with SU(2) spin symmetry, U(1) charge conservation, and particle-hole symmetry.
Although the minimal effective model for the AIM is the Kondo model Eq. 25, the MC effective model $\hat{H}^{\rm MC}_{\text{eff}}(J,U_{0})$, with $\tilde{H}^{\rm imp}_{\rm eff}$ parameterized according to Eq. 47, is precisely the Kondo model plus the leading RG-irrelevant perturbation to the local moment fixed point identified by Krishnamurthy, Wilkins and Wilson in Ref. [51]. The extended-impurity approach thus allows us to define more general models that include such RG marginal and irrelevant terms, beyond those obtained by e.g. a Schrieffer-Wolff transformation.
When the AIM is the bare model, the extended impurity Fock space dimension is reduced modestly from 16 to 8, but the parametric complexity $(U_{d},V)\to(J,U_{0})$ is the same and we do not appear to have gained much. However, we emphasize that the AIM is not the intended application of this method. Rather, we are interested in being able to systematically derive expressive effective models for complex impurity systems. We provide examples of this in the following. The AIM is however still a useful benchmark test case which we continue to explore.
III.2 UML optimization of the MC effective model
With the form of the MC effective model now at hand, we turn to the problem of optimizing its parameters using the UML method. As explained in Sec. I, this amounts to matching the observables $\langle\hat{\Omega}^{\dagger}_{\rm ad}\hat{h}_{i}^{\phantom{\dagger}}\hat{\Omega}^{\phantom{\dagger}}_{\rm ad}\rangle_{\rm bare}=\langle\hat{h}_{i}\rangle_{\rm eff}$ in bare and effective models, where $\hat{h}_{i}$ are the operators comprising the MC effective model, and $\hat{\Omega}_{\rm ad}$ is the admissibility operator that projects onto the reduced Fock space of the effective model. The explicit form of these will be discussed below, and an explicit example of the UML optimization given at the level of the AIM.
First, we note that for the MC effective model (in either form, Eq. 45 or Eq. 47), we have two observables to match and two parameters to learn when comparing with the bare model. In particular, we emphasize that the identity $\hat{\mathbbm{1}}_{\rm eff}$ is not a matchable observable, because the Fock spaces of bare and effective models are different by construction. Therefore although $\langle\hat{\mathbbm{1}}_{\rm eff}\rangle_{\rm eff}=1$ in the effective model, we have $\langle\hat{\Omega}^{\dagger}_{\rm ad}\cdot\hat{\mathbbm{1}}_{\rm eff}\cdot\hat{\Omega}_{\rm ad}\rangle_{\rm bare}<1$ in the bare model. From this it follows that it is in fact impossible to demand that,
be satisfied simultaneously for the full set of projectors $\hat{P}_{QS}$. Importantly, we conclude that the optimization problem Eq. 9 is overdetermined when using these projectors. Therefore at least one projector must be omitted from the UML observable matching process. In Eq. 45 we eliminated $\hat{P}_{0,0}$, but other choices can be made.
Another key point is that matching different combinations of observables may yield different results for the optimized effective model parameters. The operator content and scaling dimensions for different observables means they are renormalized differently. Since bare and effective models have different RG flows in the UV, different effective model parameters are in general needed to ensure agreement between for a given set of observables. We return to the systematic selection of optimal observables below; but here we simply note that observables that flow the least are the best choice.
Although the impurity Hamiltonian terms in $\hat{H}_{\rm eff}^{JU_{0}}$ in Eq. 47 are arguably more physically meaningful, the decomposition in terms of symmetry projectors in Eq. 45 turns out to be most advantageous for UML. This is because (i) the projectors have a direct meaning in the bare model, since bare and effective models have the same symmetries; (ii) the projectors mutually commute, and so the optimization problem is convex; (iii) the action of the admissibility operator on the projectors is straightforward.
III.2.1 UML for the AIM
We first look at the AIM as our bare model. As explained above, we have two observables to match within the UML optimization procedure. Working with the projectors in Eq. 45 we tune the effective model parameters $\theta_{\pm 1,1/2}$ and $\theta_{0,1}$ via GD according to Eq. 10. Importantly, the KLD loss function itself need not be computed at any stage; the optimization can be done with a knowledge of its gradient, which involves only the observables $\langle\hat{P}_{\pm 1,1/2}\rangle_{\rm eff}$ and $\langle\hat{P}_{0,1}\rangle_{\rm eff}$ in the effective model, and $\langle\hat{\Omega}^{\dagger}_{\rm ad}\hat{P}_{\pm 1,1/2}\hat{\Omega}^{\phantom{\dagger}}_{\rm ad}\rangle_{\rm bare}$ and $\langle\hat{\Omega}^{\dagger}_{\rm ad}\hat{P}_{0,1}\hat{\Omega}^{\phantom{\dagger}}_{\rm ad}\rangle_{\rm bare}$ in the bare model. The optimal solution is found when both: